Suppose a consumers preferences can be represented by the utility function U (x,y)=min(2x,y).also suppose the consumer has 300 dollar to spend and the price of good x is px =3 dollar and the price of good y is py =1 dollar.if the consumer maximize their utility subject to their budget constraint,how much of good x and how much of good y will the consumer purchase?


1
Expert's answer
2021-09-06T13:48:36-0400

Maximize U (x,y)=min(2x,y)

Subject to 3x + y ≤ 300

x≥0, y≥0

we then form a lagrangian function to help us solve the problem

L= 2xy + λ(3x+y-300)

"\frac{∂L}{∂x}" = 2y +3λ =0...............equation 1


"\frac{∂L}{∂y}" = 2x + λ =0.................equation 2


"\frac{∂L}{∂λ}" = 3x+y-300 =0.............equation 3


making λ the subject in equations 1 and 2 and then equating them

λ = "\frac{-2y}{3}"


λ= -2x

"\frac{-2y}{3}" =-2x

x = "\frac{4y}{3}"

substituting x = "\frac{4y}{3}" in equation 3

3x+y =300

"\frac{3(4y)}{3}" +y =300

5y = 300

y = 60

substituting y = 60 in x = "\frac{4y}{3}"

x = "\frac{4\times60}{3}"

x= 80

Therefore the consumer will purchase 80 units of X and 60 units of y